OpenStudy (patd):

Please help!! Solve for x: 2/5(x - 2) = 4x

OpenStudy (laura.swart):

2/5x-4/5=4x Subtract 2/5 from 4 (4 is equal to 20/5) so you get -4/5=18/5x Divide both side (flip and multiply) so -4/5 times 5/18 gives you: -72/25

OpenStudy (patd):

Thanks so much, I understand it now.